find the roots of x^2-2x=-5
Answer by edjones(7569)

x^2-2x=-5
x^2-2x+1=-5+1 Completing the square.
(x-1)^2=-4
x+1=+-sqrt(-4)
x=-1+-2i
x=-1+2i, x=-1-2i
.
Ed